# Lee-Wick instability; pion condensate; strangelets; strange stars

### v1

@@ 3.2 Dibaryons and neutral light bosons
The possibility of the existence of multiquark states was predicted by QCD-inspired models Jaffe (1977); Matveev and Sorba (1977). These works initiated a lot of experimental searches for six-quark states (dibaryons). Usually, one looked for dibaryons in the $NN$ channel. Such dibaryons would have decay widths from a few up to a hundred MeV. The obtained results were not very conclusive and the background determination was questionable. Enthusiasm has been revived in the field after experimental reports Tatischeff et al. (1997); Fil’kov et al. (2000, 2001). Reaction $pp\rightarrow p\pi^{+}\,X$ was studied in Tatischeff et al. (1997) and three narrow peaks (width $\sim 5$ MeV) in the missing mass spectrum were seen at $M_{X}=1004$ , $1044$ and $1094$ MeV with high statistical significance. The further study in Tatischeff et al. (2003) indicated the narrow baryonic structures at 1004, 1044, 1094, 1136, 1173, 1249, 1277, and 1384 MeV (and possibly 1339 MeV) in the missing mass $M_{X}$ and in the invariant mass $M_{p\pi^{+}}$. The reaction $pd\rightarrow ppX_{1}$ was studied in Fil’kov et al. (2000, 2001) and three peaks of the 5 MeV width were clearly observed in the missing mass $M_{pX_{1}}$ spectrum at $M_{pX_{1}}=1904\pm 2$ , $1926\pm 2$ and $1942\pm 2$ MeV. In the missing mass $M_{X_{1}}$ spectrum the peaks are located at $M_{X_{1}}=966\pm 2$ , $986\pm 2$ and $1002$ MeV. The peaks in the $pd$ reaction can be interpreted as the signatures of either dibaryons or nucleon resonances. Theoretical and experimental bounds on the properties of nucleon resonances were discussed in Ref. Azimov et al. (2003). The resonance states below the pion-nucleon threshold and their consequences for nuclear systems were discussed in Kolomeitsev and Voskresensky (2003).
@@
Rather recently several very heavy compact stars were observed. The current highest precisely measured mass is $2.08(7)\,M_{\odot}$ for PSR J0740+6620 Cromartie et al. (2019); Fonseca et al. (2021). On the other hand appearance of hyperons and $\Delta$ resonances in neutron star interiors results in a decrease of the maximal neutron star mass. Nevertheless these so called the hyperon and $\Delta$ isobar puzzles can be resolved, see Kolomeitsev et al. (2017) and reference therein. Reference Kolomeitsev and Voskresensky (2003) demonstrated that in presence of light dibaryons the maximal possible mass of a neutron star would decrease below the observational limit. Appearance of rather heavy dibaryons and baryon resonances although results in a decrease of the maximal neutron star mass does not contradict to observations.
@@
An interesting alternative interpretation of the tower of resonance states was given in Walcher (2001). It was noted that drawing all new states (both dibaryonic and nucleonic ones) on one energy scale, one obtains a tower of almost equidistant states. It was hypothesized the existence of a light pseudoscalar meson ($J^{P}=0^{-}$) with a mass of $m_{{\rm light}\,\pi}\simeq 21\pm 2.6$ MeV. The basic idea is that the tower of excited nucleonic states is formed by the nucleon in its ground state plus $1,2,3,\dots$ light pions as the quantum of excitation with the energy $m_{{\rm light}\pi}$. It was argued in Kolomeitsev and Voskresensky (2003) that the presence of light charged pions would allow for existence of abnormal nuclei ($A\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10^{3}$) and “nuclei stars” of arbitrary size, bound by strong and electromagnetic interactions.
@@
Recent possible discovery is also worth mentioning: a light neutral boson (X17) with a mass of 17 MeV and a very narrow width (presumably less than $0.07$ eV) was found at the accelerator of ATOMKI. Three significant anomalies were observed in electron–positron pairs emitted in the 3H $(p,e^{-}e^{+})$ 4He, 7Li $(p,e^{-}e^{+})$ 8Be, and 11B $(p,e^{-}e^{+})^{12}$C reactions Krasznahorkay et al. (2016, 2021, 2022). If confirmed, the existence of this new particle could be important for particle physics and cosmology, see review Gustavino (2024). Reference Feng et al. (2016) assumed that X17 is the vector gauge boson coupled more strongly to neutrons than to protons. Reference Ellwanger and Moretti (2016) suggested another interpretation of the experimental results, assuming a light, pseudoscalar particle. In both cases the coupling of X17 with nucleons was assumed to be very weak.
@@
On the other hand, independently of the value of the strength of the interaction it is worthwhile to ask, if existence of a light neutral boson does not contradict to the neutron star properties. Reference Kanakis-Pegios et al. (2024) assumed that X17 can mediate low-energy nucleon-nucleon interactions and attempted to find possible constraints on the hypothetical X17 boson from the analysis of the equation of state of neutron stars. It was shown that such a neutron star could have MRM-R curve rather similar to that for the ordinary neutron star.
@@
Note that since $m_{X17}/m_{\pi}\sim 1/8$ one could expect presence of the nucleon–light boson polarization effects of the same order as for pions, if the coupling of the light boson with nucleons were $f_{XNN}\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}f_{\pi NN}/8$. In the latter case a possibility would exist for the Bose condensation of such a particles in the compact star matter (similarly to $\pi^{0}$ condensation) and feasibly for existence of metastable or stable nuggets and light-boson enriched nuclei-stars. Nevertheless we should repeat that Ellwanger and Moretti (2016) and Feng et al. (2016) estimated X17-nucleon polarization effects are tiny.
3.3 Rotating nuclear systems
@@
The fastest known pulsar, PSR J1748-2446ad, makes one revolution in 1.4 ms (angular frequency $\Omega\approx 4.5\times 10^{3}$ Hz). Thus typical angular velocities of so far observed neutron stars are $\Omega R\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}{0.15c}$, where $c$ is the speed of light and $R\sim 10$ km is the neutron star radius. For a compact star rotating at the frequency close to the Kepler frequency $\Omega_{G}\sim\sqrt{GM/R^{3}}$, the rotation affects the equation of state resulting in an increase of the mass of the static star on the value $\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}(0.3-0.4)M_{\odot}$, cf. Farrell et al. (2024). When such a neutron star will undergo collapse to a black hole, its angular rotation frequency will increase due to conservation of the angular momentum up to $\Omega<1/r_{G}=1/(2MG)\sim 10^{5}$Hz for the compact star with the mass $M\sim M_{\odot}$, where $M_{\odot}\approx 2\times 10^{33}$ g is the solar mass, and $G$ is the gravitational constant. For primordial black holes with masses $M<10^{14}$ g, this estimate would give $\Omega\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}10^{24}$ Hz.
@@
Condensation of pion excitations on a low-lying branch of the excitation spectrum in rectilinearly moving and rotating media was considered in Voskresensky (1993). Previously it was thought that a compact star becomes unstable provided the rotation frequency is enlarger than the Kepler value $\Omega>\Omega_{G}$. However Ref. Voskresensky (1993) argued that a part of the angular momentum can be transported to the forming condensate of Bose excitations, and as the result the angular frequency of the normal component may decrease below $\Omega_{G}$.
Magnetic fields in ordinary pulsars, like the Crab pulsar, reach values $\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}10^{13}$ G at their surfaces. At the surface of magnetars magnetic fields are higher, $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10^{15}$ G. In the interior the magnetic field might be even stronger (up to $\sim 10^{18}$ G) depending on the assumed mechanism of the formation of the magnetic field Voskresensky and Anisimov (1980). Also, magnetic fields $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}(10^{17}\mbox{--}10^{18})$ G, $h\sim h_{\rm VA}=Ze/R^{2}\sim H_{\pi}(Ze^{6})^{1/3}$ (where typical time of the collision process was estimated as $\Delta t\sim R/c$) may exist in non-central heavy ion collisions at collision energies $\sim$ GeV/A, as it was evaluated in Ref. Voskresensky and Anisimov (1980), $Z$ is the charge of the fireball, $H_{\pi}=m_{\pi}^{2}/e\simeq 3.5\times 10^{18}$ G.
In presence of a magnetic field and electrical potential well there exists the critical angular velocity, above which from the vacuum in the rotation frame there arises the pion condensate. In the frame rotating with the angular velocity $\vec{\Omega}\parallel z$ the solution for the charged condensate field has the form of the vortex, $e^{il\theta}\phi$, where $\theta$ is the angle in the cylindrical system and $l$ is the quantum number of the angular momentum. To be specific let further $\Omega,l>0$. Then equations for the boson field and the electric potential (18) and (19) are to be replaced by the following ones Voskresensky (2024, 2025),
@@
where $A_{\theta}(r)$ is the angular component of the vector potential of the magnetic field, and a $\phi$-field self-interaction is for simplicity neglected. Equations (21), (22) should be supported by equation for $\vec{A}$. However with the ansatz $A_{\theta}(r))=l/r$ we have $\mbox{curl}\vec{A}=0$ and equation for $\vec{A}$ is fulfilled identically.
@@
Now in case of a nuclear system ($N\simeq Z$, $n\approx n_{0}$) of a large size solutions are $V=\Omega l-m_{b}\leq 0$, $|\phi|^{2}=n_{p}/(2m_{b})$ and instead of (20) we find
$\mathcal{E}=-\mathcal{E}_{0}A+\frac{1}{2}(m_{b}-\Omega l)A+\mathcal{E}_{kin}\,$, (23)
@@
where ${\cal{E}}_{kin}$ is the kinetic energy of the rotation of the nucleon subsystem. The term $lA/2$ is smaller or equal to the initial angular momentum of the charged uniformly rotating nucleon sub-system $L\sim Am_{N}\Omega R^{2}$ (before appearance of the pion giant vortex, $l\gg 1$), $m_{N}$ is the mass of the nucleon. Thus $\Omega$ is determined by the initial value of the total angular momentum. Then $l\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}m_{N}\Omega R^{2}$. The critical value of $l$ is determined by the condition $\Omega l_{c}=m_{b}-2\mathcal{E}_{0}$. In the rotating frame the nuclearite remains bound till the system will not be decelerated by the weak radiation force. For a large system the latter time can be very long.
@@
The rotation decreases the value of the effective pion mass, see (21), thereby favoring the pion condensation at a smaller critical density. This observation motivates a possibility of presence of the differential rotation characterizing by a more rapidly rotating charged pion superfluid (favoring a stronger condensate) and a more slowly rotating surrounding nucleon exterior. Notice that differentially rotating stars can support more mass than their uniformly rotating or non-rotating counterparts Morrison et al. (2004). In the case of massive remnant stars of binary merger events, differential rotation is one of the mechanisms that may provide an extra centrifugal support to stabilize the star well above the Tolman-Oppenheimer-Volkoff (uniform rotation) mass limit.
3.4 Anomalous supercharged nuclei consisting dark matter

### v2

@@ III.2 Dibaryons and neutral light bosons
The possibility of the existence of multiquark states was predicted by QCD-inspired models [135, 199]. These works initiated a lot of experimental searches for six-quark states (dibaryons). Usually, one looked for dibaryons in the $NN$ channel. Such dibaryons would have decay widths from a few up to a hundred MeV. The obtained results were not very conclusive and the background determination was questionable. Enthusiasm has been revived in the field after experimental reports [266, 84, 83]. Reaction $pp\rightarrow p\pi^{+}\,X$ was studied in [266] and three narrow peaks (width $\sim 5$ MeV) in the missing mass spectrum were seen at $M_{X}=1004$ , $1044$ and $1094$ MeV with high statistical significance. The further study in [267] indicated the narrow baryonic structures at 1004, 1044, 1094, 1136, 1173, 1249, 1277, and 1384 MeV (and possibly 1339 MeV) in the missing mass $M_{X}$ and in the invariant mass $M_{p\pi^{+}}$. The reaction $pd\rightarrow ppX_{1}$ was studied in [84, 83] and three peaks of the 5 MeV width were clearly observed in the missing mass $M_{pX_{1}}$ spectrum at $M_{pX_{1}}=1904\pm 2$ , $1926\pm 2$ and $1942\pm 2$ MeV. In the missing mass $M_{X_{1}}$ spectrum the peaks are located at $M_{X_{1}}=966\pm 2$ , $986\pm 2$ and $1002$ MeV. The peaks in the $pd$ reaction can be interpreted as the signatures of either dibaryons or nucleon resonances. Theoretical and experimental bounds on the properties of nucleon resonances were discussed in Ref. [16]. The resonance states below the pion-nucleon threshold and their consequences for nuclear systems were discussed in [163].
@@
Rather recently several very heavy compact stars were observed. The current highest precisely measured mass is $2.08(7)\,M_{\odot}$ for PSR J0740+6620 [56, 87]. On the other hand appearance of hyperons and $\Delta$ resonances in neutron star interiors results in a decrease of the maximal neutron star mass. Nevertheless these so called the hyperon and $\Delta$ isobar puzzles can be resolved, see [152] and reference therein. Reference [163] demonstrated that in presence of light dibaryons the maximal possible mass of a neutron star would decrease below the observational limit. Appearance of rather heavy dibaryons and baryon resonances although results in a decrease of the maximal neutron star mass does not contradict to observations.
@@
An interesting alternative interpretation of the tower of resonance states was given in [293]. It was noted that drawing all new states (both dibaryonic and nucleonic ones) on one energy scale, one obtains a tower of almost equidistant states. It was hypothesized the existence of a light pseudoscalar meson ($J^{P}=0^{-}$) with a mass of $m_{{\rm light}\,\pi}\simeq 21\pm 2.6$ MeV. The basic idea is that the tower of excited nucleonic states is formed by the nucleon in its ground state plus $1,2,3,\dots$ light pions as the quantum of excitation with the energy $m_{{\rm light}\pi}$. It was argued in [163] that the presence of light charged pions would allow for existence of abnormal nuclei ($A\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10^{3}$) and “nuclei stars” of arbitrary size, bound by strong and electromagnetic interactions.
@@
Recent possible discovery is also worth mentioning: a light neutral boson (X17) with a mass of 17 MeV and a very narrow width (presumably less than $0.07$ eV) was found at the accelerator of ATOMKI. Three significant anomalies were observed in electron–positron pairs emitted in the 3H $(p,e^{-}e^{+})$ 4He, 7Li $(p,e^{-}e^{+})$ 8Be, and 11B $(p,e^{-}e^{+})^{12}$C reactions [165, 164, 166]. If confirmed, the existence of this new particle could be important for particle physics and cosmology, see review [114]. Reference [82] assumed that X17 is the vector gauge boson coupled more strongly to neutrons than to protons. Reference [76] suggested another interpretation of the experimental results, assuming a light, pseudoscalar particle. In both cases the coupling of X17 with nucleons was assumed to be very weak.
@@
On the other hand, independently of the value of the strength of the interaction it is worthwhile to ask, if existence of a light neutral boson does not contradict to the neutron star properties. Reference [140] assumed that X17 can mediate low-energy nucleon-nucleon interactions and attempted to find possible constraints on the hypothetical X17 boson from the analysis of the equation of state of neutron stars. It was shown that such a neutron star could have MRM-R curve rather similar to that for the ordinary neutron star.
@@
Note that since $m_{\rm X17}/m_{\pi}\sim 1/8$, one could expect presence of the nucleon–light boson polarization effects of the same order as for pions, if the coupling of the light boson with nucleons were $f_{XNN}\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}f_{\pi NN}/8$. In the latter case a possibility would exist for the Bose condensation of such a particles in the compact star matter (similarly to $\pi^{0}$ condensation) and feasibly for existence of metastable or stable nuggets and light-boson enriched nuclei-stars. Nevertheless we should repeat that [76] and [82] estimated X17-nucleon polarization effects as tiny.
III.3 Rotating nuclear systems
@@
The fastest known pulsar, PSR J1748-2446ad, makes one revolution in 1.4 ms (angular frequency $\Omega\approx 4.5\times 10^{3}$ Hz). Thus typical angular velocities of so far observed neutron stars are $\Omega R\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}{0.15c}$, where $c$ is the speed of light and $R\sim 10$ km is the neutron star radius. The limiting rotation frequency of a compact star — the Kepler frequency — is determined by equation $m_{N}\Omega_{\rm K}^{2}R^{2}/2=GMm_{N}/R$, where $G$ is the gravitational constant, $R$ is the radius of the object, and $M$ is its mass. For a compact star rotating with the frequency close to the Kepler frequency $\Omega_{\rm K}$, the rotation affects the equation of state resulting in an increase of the mass of the static star on the value $\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}(0.3-0.4)M_{\odot}$, cf. [81]. When such a compact star collapses to form a black hole, its angular rotation frequency increases due to conservation of the angular momentum up to $\Omega<1/r_{G}=1/(2MG)\sim 10^{5}$Hz for the compact star with the mass $M\sim M_{\odot}$, where $M_{\odot}\approx 2\times 10^{33}$ g is the solar mass. For primordial black holes with masses $M<10^{14}$ g, this estimate would give $\Omega\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}10^{24}$ Hz.
@@
Magnetic fields in ordinary pulsars, like the Crab pulsar, reach values $\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}10^{13}$ G at their surfaces. At the surface of magnetars magnetic fields are higher, $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10^{15}$ G. In the interior the magnetic field might be even stronger (up to $\sim 10^{18}$ G) depending on the assumed mechanism of the formation of the magnetic field [274]. Also, magnetic fields $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}(10^{17}\mbox{--}10^{18})$ G, $h\sim h_{\rm VA}=Ze/R^{2}\sim H_{\pi}(Ze^{6})^{1/3}$ (where typical time of the collision process was estimated as $\Delta t\sim R/c$) may exist in non-central heavy ion collisions at collision energies $\sim$ GeV/A, as it was evaluated in Ref. [274], $Z$ is the charge of the fireball, $H_{\pi}=m_{\pi}^{2}/e\simeq 3.5\times 10^{18}$ G.
As in Sect. III.1 we consider now negatively charged scalar bosons (or pseudoscalar pions) with a mass $m_{b}$. In a presence of a magnetic field and an electrical potential well, there exists a critical angular velocity, above which, from the vacuum in the rotation frame there arises a condensate of the charged bosons [187, 113, 289, 291, 38]. In the frame rotating with the angular velocity $\vec{\Omega}\parallel z$, the solution for the charged condensate field has the form of a vortex, $e^{il\theta}\phi$, where $\theta$ is the angle in the cylindrical system and $l$ is the quantum number of the angular momentum. To be specific let further $\Omega$ and $l$ be positive. In the presence of a uniform constant magnetic field $\vec{H}\parallel z$ and a rectangular potential well $V=-V_{0}=const$ for $\sqrt{x^{2}+y^{2}}<R$, the critical angular velocity is given by $l\Omega_{c}=-V_{0}+\sqrt{m^{2}_{b}+|eH|(1-2\alpha)}$ with $1\leq\alpha\leq 0$.
For the case of the rotating (for simplicity cylindric) nucleus equations for the boson field and the electric potential (18) and (19) are to be replaced by the following ones [289, 291],
@@
where $A_{\theta}(r)$ is the angular component of the vector potential of the magnetic field, and a $\phi$-field self-interaction is for simplicity neglected. Equations (21), (22) should be supported by equation for the vector-potential $\vec{A}$. However with the ansatz that $A_{\theta}(r)=l/r$ in the vortex exterior we have $\mbox{curl}\vec{A}=0$ and equation for $\vec{A}$ is fulfilled identically.
@@
Now in case of a nuclear system ($N\simeq Z$, $n\approx n_{0}$) of a large size solutions are $V=\Omega l-m_{b}\leq 0$, $|\phi|^{2}=n_{p}/(2m_{b})$ and instead of (20) we find  [289, 291]
$\mathcal{E}=-\mathcal{E}_{0}A+\frac{1}{2}(m_{b}-\Omega l)A+\mathcal{E}_{kin}\,,\,A=N+Z,$ (23)
@@
where ${\cal{E}}_{kin}$ is the kinetic energy of the rotation of the nucleon subsystem. The term $lA/2$ is smaller or equal to the initial angular momentum of the charged uniformly rotating nucleon sub-system $L\sim Am_{N}\Omega R^{2}$ (before appearance of the pion giant vortex, $l\gg 1$), $m_{N}$ is the mass of the nucleon. At the rigid rotation $\Omega$ is determined by the initial value of the total angular momentum. The critical values of $l$ and $\Omega$ are determined by the condition $\Omega_{c}l_{c}=m_{b}-2\mathcal{E}_{0}$. The role of $m_{b}$ can be played by the effective mass of a charged pion, which decreases with an increase of the baryon density, see Sect. IV below. Here we considered the possibility of the second-order phase transition to a state with a giant vortex of the charged pion field. In reality, the phase transition can be of the first order with a jump of the condensate field at the critical point. Then the critical angular velocity is given by the condition $\Omega_{c}l_{c}=m_{b}-2\mathcal{E}_{0}+2\mathcal{E}_{\rm cond}(\Omega_{c}l_{c})$, where $\mathcal{E}_{\rm cond}\leq 0$ is a gain in the energy per baryon due to the formation of the vortex. In the rotating frame the nugget remains self-bound until the system is not decelerated by the weak radiation force below $\Omega_{c}$. For a large-size system the deceleration time can be very long.
Now we mention another possibility. In the case of rectilinearly moving and rotating media, condensation of Bose excitations on a low-lying branch of the excitation spectrum $\omega(k)$, behaving similar to that in superfluid 4He and in Fermi liquids, was studied in Ref. [283], see also Ref. [159] and references therein. As shown, when the speed of rectilinear motion exceeds a critical value (typically Fermi velocity for a nucleon medium) a part of the momentum can be transferred from the normal nucleon system to the superfluid condensate of Bose excitations with energy $\omega(k_{c})$. Similarly when the rotation frequency exceeds some critical value a part of the angular momentum can be transferred from the normal nucleon system to the superfluid condensate of Bose excitations [283, 157]. Consequently, the angular frequency of the normal component may decrease below $\Omega_{\rm K}$, whereas it would exceed $\Omega_{\rm K}$ otherwise.
When the baryon density in the interior of the star exceeds the value $n_{c}^{\pi^{+}}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}n_{0}$, there appears a new branch of $\pi^{+}$ excitations with a negative frequency $\omega(k)$ and it may become energetically profitable to transport a part of the angular momentum to the $\pi^{+}$ condensate droplets, for a more detail see Ref. [283] and Sect. IV.2 below. This allows to propose the following scenario. When during the initial compression stage of a supernova collapse, the density exceeds the critical density $n_{c}^{\pi^{+}}$ it becomes energetically favorable to transfer a part of the initial angular momentum from the nucleon subsystem to pion condensate drops. If the star is destroyed in the end, these drops will continue to rotate (with angular velocities $\Omega R_{\rm d}<1$, with $R_{\rm d}$ being a size of the drop) until the angular momentum has been completely radiated from the drop surface. These drops of $\pi$ condensate nuclear matter may be observed as abnormal nuclei and nuclearites propagating in cosmos.
@@
Notice also that differentially rotating stars can support a larger star mass than their uniformly rotating or non-rotating counterparts [216]. In the case of massive remnant stars of binary merger events, differential rotation is one of the mechanisms that may provide an extra centrifugal support to stabilize the star well above the Tolman-Oppenheimer-Volkoff (uniform rotation) mass limit.
III.4 Anomalous supercharged nuclei consisting dark matter